# Split & Merge: Unlocking the Potential of Visual Adapters via Sparse Training

### 05 Dec 2023 v1

@@ Abstract
With the rapid growth in the scale of pre-trained foundation models, parameter-efficient fine-tuning techniques have gained significant attention, among which Adapter Tuning is the most widely used. Despite achieving efficiency, Adapter Tuning still underperforms full fine-tuning, and the performance improves at the cost of an increase in parameters. Recent efforts address this issue by pruning the original adapters, but it also introduces training instability and suboptimal performance on certain datasets. Motivated by this, we propose Mixture of Sparse Adapters, or MoSA, as a novel Adapter Tuning method to fully unleash the potential of each parameter in the adapter. We first split the standard adapter into multiple non-overlapping modules, then stochastically activate modules for sparse training, and finally merge them to form a complete adapter after tuning. In this way, MoSA can achieve significantly better performance than standard adapters without any additional computational or storage overhead. Furthermore, we propose a hierarchical sparse strategy to better leverage limited training data. Extensive experiments on a series of 27 visual tasks demonstrate that MoSA consistently outperforms other Adapter Tuning methods as well as other baselines by a significant margin. Furthermore, in two challenging scenarios with low-resource and multi-task settings, MoSA achieves satisfactory results, further demonstrating the effectiveness of our design. Our code will be released. footnotemark: footnotetext: Corresponding author. Code: https://github.com/Theia-4869/MoSA.
@@ 1 Introduction
Figure 1: Different Adapter Tuning diagrams: (a) Adapter Tuning schema and standard adapter. (b) Sparse adapter. (c) Our proposed Mixture of Sparse Adapters (MoSA). We first split the standard adapter into multiple non-overlapping modules, then stochastically activate them for sparse training, and finally merge them to form a complete dense adapter.
@@
In computer vision, the pre-training paradigm has witnessed remarkable success [7, 22, 23]. Models pre-trained on large-scale datasets (e.g. ImageNet-21k [60], JFT-300M [62], SA-1B [32]) have demonstrated significant performance improvements across various downstream tasks. With pre-trained models, transfer learning aims to transfer the acquired knowledge to novel target domains, where labeled training samples are typically insufficient [53, 3, 79]. One of the most direct methods is fine-tuning, wherein pre-trained model parameters are further updated based on downstream task data [71, 33]. Currently, the widely adopted fine-tuning approach is full fine-tuning, involving the update of all parameters in the pre-trained model during tuning. However, as the scale of pre-trained models continues to escalate (e.g. ViT-G/14 [76] with over 1.8B parameters), storing the complete set of parameters for each task becomes impractical, giving rise to the field of parameter-efficient fine-tuning (PEFT), which addresses this challenge by updating only a small subset of parameters from the pre-trained model for each downstream task. PEFT not only ensures the efficiency of parameter storage but also optimizes the utilization of available data.
Recently, various PEFT methods have emerged [74, 38, 37, 41, 29, 27, 39], with one of the most notable being Adapter Tuning [26, 57, 58, 6, 8, 52, 54]. This kind of method introduces additional lightweight modules to the pre-trained model while freezing the model backbone, achieving impressive results. Adapters [26, 57, 58] first make advancements in natural language processing. AdaptFormer [6], on the other hand, changes the adapter insertion form from sequential to parallel, achieving better results in ViT [13] architecture and visual tasks. Moreover, there is a growing trend in the visual domain towards adapters designed specifically for visual tasks [52, 54]. Despite the consideration for parameter efficiency, these adapters still underperform standard full fine-tuning. Increasing the bottleneck dimension in adapters can raise performance, but this comes at the cost of introducing too many additional parameters. Recent efforts have explored pruning the entire model [19, 70, 1] or adapter [24] before tuning to enhance the efficiency of parameter utilization, seemingly mitigating this problem to some extent. However, excessive sparsity can lead to instability and suboptimal performance on certain datasets. The delicate balance between parameter efficiency and performance remains a key challenge in Adapter Tuning methods.
@@
In this paper, we propose Mixture of Sparse Adapters (MoSA) as a solution to the aforementioned issues, enhancing adapter performance without introducing additional computational and storage overhead. This is achieved through the incorporation of sparsity and mixture-of-expert (MoE) [61, 59, 78]. We start by splitting the standard adapter into several non-overlapping modules, each can be considered as a sparse adapter. Since the parameters of each module need to be updated, we employ a random way for decomposition, and experiments show that this approach is both concise and effective. During the training phase, to avoid extra computational costs, we use a non-routing stochastic activation mechanism. Each activated module uses a mask to filter gradients, achieving sparsity in parameter updates. We merge all sparse adapters into a complete dense adapter, achieving efficiency in both storage and computation during inference. Through this design, the potential of the original adapter structure is fully unleashed, maximizing the elimination of parameter redundancy.
@@
Rather than simply stacking two independent approaches, our design organically integrates the sparse training with MoE, where the two components mutually reinforce each other. On one hand, the training process with mixed experts ensures expressive capability and stability in downstream tasks, addressing the limitations of sparse training. On the other hand, sparse update of parameters more efficiently utilizes training data, alleviating the data dilution caused by sparse activation of multiple experts. To better facilitate the combination of the two components, we further propose a hierarchical sparse strategy. The dense down-projection layer provides robust intermediate features for the model, while multiple sparse up-projection layers increase the capacity of the model. It’s worth noting that our approach, as a general concept of enhancing parameter utilization, can be applied to various adapter structures and other addition-based PEFT methods. On 27 diverse downstream visual tasks with a pre-trained ViT backbone, our MoSA consistently outperforms all other fine-tuning methods, including full fine-tuning. Compared to full fine-tuning, MoSA achieves a lead of 1.36% (on FGVC), 2.43% (on VTAB) and 1.51% (on GICD), while updating only around 1% of the parameters compared to the pre-trained backbone. Compared to the standard adapter, MoSA achieves an improvement of 1.32% (on FGVC), 1.06% (on VTAB) and 0.36% (on GICD) without adding any computational or storage overhead. Additionally, we conducted comprehensive ablation experiments to validate the effectiveness of each component in our design.
@@
  • We propose a novel Adapter Tuning method, namely MoSA, for fully unleashing the potential of the standard adapters. Through a sparse and mixed training approach involving splitting and merging, our method maximizes parameter efficiency, enhancing the performance of Adapter Tuning in visual tasks.
  • MoSA maximizes the balance between efficiency and performance. It exhibits efficiency in all the sparsification, mixed training, and inference, and the organic integration of sparsity and MoE further enhances performance.
  • We evaluate our method on a total of 27 downstream visual tasks spanning different domains, and MoSA significantly outperforms full fine-tuning as well as all other PEFT baselines. In addition to common scenarios, MoSA also achieves optimal results in challenging scenarios such as low-resource and multi-task settings, further demonstrating the rationality of our design.
@@ 2 Related work
2.1 Parameter-efficient fine-tuning (PEFT)
Recently, many large-scale pre-trained models (e.g. LLaMA2 [63] 70B, GPT-3 [5] 175B) have emerged in deep learning research, which can achieve excellent performance in a variety of downstream tasks. However, updating and storing all model parameters for each downstream task has become far more expensive. PEFT achieves efficiency in training and storage by updating only a small fraction of parameters compared to pre-trained models [74, 38, 37, 29, 27, 30, 39, 40, 42], among which Adapter Tuning is one of the most widely used methods. Adapter [26] updates only the parameters of newly added lightweight bottleneck modules while freezing the model backbone during fine-tuning, achieving great success in NLP tasks. AdapterFusion [57] adjusts the position of the adapter and combines the parameters of adapters in multiple tasks to better adapt to multi-task scenarios. AdaptFormer [6] improves the performance of ViT [13] on visual tasks by changing the connection pattern of adapters from sequential to parallel. Moreover, ST-Adapter [52] and DualPath [54] use adapters to do image-to-video transfer learning. In contrast to the methods mentioned above, in this paper, we propose a new framework for training multiple sparse adapters, aiming to enhance performance while maintaining the efficiency of the original Adapter Tuning approach.
2.2 Mixture-of-experts (MoE)
@@
The origin proposal of MoE [28] is targeted at enhancing model capacity. The earliest proposed technologies, known as soft MoE [47], integrate outputs of multiple experts through weighted summation. However, this technique significantly increases the computational cost during training. To address this issue, the sparsely-gated MoE [61, 78, 21] was introduced, selecting specific experts for activation and directly assigning data to them, thereby reducing the computational burden. Nonetheless, this method often resulted in load imbalance, with some experts becoming inactive in later training stages, risking system collapse. THOR [80] effectively mitigates this issue by implementing random activation, which not only enhances efficiency but also improves overall model performance. Unlike most previous methods in the fields, which typically employ the complete model as a singular expert encompassing all functionalities, our approach innovatively utilizes the sparse PEFT module as a distinct expert within the training process. After completing the training phase, we strategically combine these modules to form a cohesive and optimized system, which is pivotal for maximizing parameter efficiency, ensuring that our model not only maintains high performance but also operates with reduced computational demand.
2.3 Pruning and sparse training
@@
The primary goal of model pruning [20, 50, 46] originally aimed to minimize deployment costs by reducing model parameters without significantly impacting its performance. Recent studies [19, 70, 1] have revealed that pruning can enhance model fine-tuning. A reduced number of trainable parameters can act as an additional regularization constraint, potentially improving performance. As a PEFT method, the architecture of adapters [26, 57] inherently possesses few updatable parameters. SparseAdapter [1] extends this concept by pre-pruning the Adapter, enhancing parameter efficiency while maintaining or even improving performance compared to standard adapters. However, excessive sparsity can lead to training instability and suboptimal results on certain datasets. We take one step further to solve aforementioned weakness via integrating the MoE paradigm into our sparse tuning process. This integration not only achieves a sparse network structure, but also implements sparse activation during the training phase, which overcomes the inherent constraints of using a single sparse network.
@@ 3 Method
Figure 2: Architecture design of Mixture of Sparse Adapters. In the training phase, MoSA stochastically activates a sparse adapter during each forward pass; in the inference phase, MoSA merges multiple sparse adapters into a complete one to enhance inference efficiency.
The adapter is a bottleneck form of PEFT module, where the standard adapter can enhance model capacity by increasing bottleneck dimensions at the cost of too many additional parameters. [24] achieves parameter efficiency through pre-pruning, but excessive sparsity can lead to instability and performance degradation. To achieve the optimal balance between efficiency and performance, we propose MoSA, which achieves efficiency through splitting and sparse training, while ensuring performance through stochastic activation and merging. To this end, we first overview our method in Sec. 3.1. Subsequently, we describe how we split the standard adapter into multiple sparse adapters in Sec. 3.2. Then, we introduce the training strategy with stochastic activation of multiple sparse experts in Sec. 3.3. Finally, in Sec. 3.4, we merge the split sparse adapters into a complete one for efficiency during inference.
@@ 3.1 Overview
where $x\in\mathbb{R}^{d}$ represents the input of the adapter, $f$ is the activation function, and $\mathcal{W}^{\text{down}}\in\mathbb{R}^{d\times r}$ and $\mathcal{W}^{\text{up}}\in\mathbb{R}^{r\times d}$ denote the linear layers for down-projection and up-projection, respectively. $r$ is the bottleneck dimension, satisfying $r\ll d$, which allows for a reduction in the number of parameters in the adapter. [24] further exploits the high efficiency of parameters by pruning $\mathcal{W}^{\text{down}}$ and $\mathcal{W}^{\text{up}}$ before tuning. In our work, we split the standard adapter into $N$ sparse adapters $\mathscr{W}^{\text{down}}=\{\mathcal{W}_{i}^{\text{down}}\}_{i=1}^{N},\mathscr{W}^{\text{up}}=\{\mathcal{W}_{i}^{\text{up}}\}_{i=1}^{N}$, and stochastically activate one of them during the training process for adaptation:
@@
where $i,j\in\{1,\cdots,N\}$, representing the stochastic selection of one from $N$ sparse adapters. The overall design of our method is shown in Fig. 2.
@@ 3.2 Sparse adapter splitting
To make the most of all the parameters within a single adapter, we first split the standard adapter into multiple sparse adapters, as illustrated in Fig. 1. As all parameters in the adapter need to be updated, unlike [24], we do not adopt the task-specific pruning mechanism but instead employ a random splitting method to achieve as balanced grouping as possible, which also avoids the additional overhead of gradient computation based on downstream tasks before fine-tuning. Given the parameter of adapter $\mathcal{W}\in\mathbb{R}^{d\times r}$, we first generate an initial score $\mathcal{S}\sim\text{Uniform}(0,1)\in\mathbb{R}^{d\times r}$ for it. Then, we split $W$ into $N$ non-overlapping sparse adapters $\mathcal{W}_{i}:i\in\{1,\cdots,N\}$. Denoting the $i$-th $N$-quantile of all values in the matrix $\mathcal{S}$ as $s_{i}:i\in\{1,\cdots,N\}$, we obtain $N$ sparse masks $\mathcal{M}_{i}$ as follows:
@@
where $\varepsilon$ represents the update step size, and $\nabla\mathcal{L}(\mathcal{W}_{i})$ denotes the gradient of the task loss with respect to $\mathcal{W}_{i}$. In this way, the positions with a mask value of 0 have their gradients filtered to 0, thereby freezing the corresponding parameters and achieving sparse training for the adapter.

### 23 Mar 2024 v2

@@ Abstract
With the rapid growth in the scale of pre-trained foundation models, parameter-efficient fine-tuning techniques have gained significant attention, among which Adapter Tuning is the most widely used. Despite achieving efficiency, it still underperforms full fine-tuning, and the performance improves at the cost of an increase in parameters. Recent efforts have either focused on training multiple adapter experts to increase model capacity or on pruning adapters to achieve parameter efficiency. However, both approaches introduce more parameters compared to the original adapter, hence are not computationally efficient. Motivated by this, we propose Mixture of Sparse Adapters, or MoSA, as a novel Adapter Tuning method to fully unleash the potential of each parameter in the adapter. We first split the standard adapter into multiple non-overlapping modules, then stochastically activate them for sparse training, and finally merge them to form a complete adapter after tuning. In this way, MoSA can achieve significantly better performance than standard adapters without any additional computational or storage overhead. Furthermore, we propose a hierarchical sparse strategy to better leverage limited training data. Extensive experiments on a series of 27 visual tasks demonstrate that MoSA consistently outperforms other Adapter Tuning methods as well as other baselines by a large margin. Furthermore, MoSA brings consistent improvements across various model scales, architectures, and different PEFT methods. Code will be released.
Keywords:
Adapter Tuning Mixture-of-Experts Sparse Training
@@ 1 Introduction
Figure 1: Different Adapter Tuning diagrams: (a) The standard adapter simply inserts a bottleneck module into each Transformer layer. (b) Sparse Adapter. It prunes the standard adapter before tuning, updating only a small subset of the retained parameters. (c) Our proposed Mixture of Sparse Adapters (MoSA). We first split the standard adapter into multiple non-overlapping modules, then stochastically activate them for sparse training, and finally merge them to form a complete dense adapter.
@@
The pretrain-then-finetune paradigm has achieved remarkable success in deep learning. Within the field of computer vision, models pre-trained on large-scale datasets (e.g. ImageNet-21k [66], JFT-300M [68], SA-1B [40]) have demonstrated significant performance improvements across various downstream tasks [9, 29, 28]. After pre-training, models require fine-tuning on specific data to transfer learned knowledge to the target domain. The most direct method is full fine-tuning, involving the update of all parameters in the pre-trained model during tuning. However, as the scale of pre-trained models continues to grow (e.g. ViT-G/14 [81] 1.8B, LVM [2] 3B), storing a complete copy of all parameters for each task becomes impractical, giving rise to a more efficient method of tuning, known as parameter-efficient fine-tuning (PEFT). For each downstream task, PEFT updates only a small portion of the the parameters in the pre-trained model, achieving efficiency in parameter storage and data utilization.
Recently, various PEFT methods have emerged [80, 46, 45, 49, 36, 39, 48, 34, 47, 85, 31], with one of the most widely used being Adapter Tuning [33, 64, 65, 7, 8, 60, 61]. This kind of method introduces lightweight bottleneck modules to the pre-trained model while freezing the backbone, as shown in Figure 1(a), facilitating efficient knowledge transfer to downstream tasks. Despite the respectable performance of these methods, they still fall behind full fine-tuning in many scenarios [25] (e.g. training data is relatively abundant, the distribution gap between the downstream and pre-training data is significant). Simply increasing the bottleneck dimension can raise performance, but this contradicts the original design philosophy of Adapter Tuning. Recent work has enhanced adapter performance by incorporating a Mixture-of-Experts (MoE) mechanism [79, 21, 16, 18], but having multiple parallel adapter modules and routers introduces additional parameters and computation. Sparsely or stochastically activated MoE is more efficient in implementation [18, 6, 73], but this reduces the amount of data seen by individual experts, an issue we call data dilution, leading to even worse performance, especially when training data is limited. Meanwhile, another way for improving adapters involves pruning before tuning [30] (e.g. a 4×4$\times$4 × larger adapter with a 75% sparse ratio), as in Figure 1(b), which has shown performance gains but does not offer computational efficiency, and high levels of sparsity can further lead to training instability. The delicate balance between parameter efficiency and performance remains a key challenge in Adapter Tuning.
@@
Therefore, a natural question arises: is it possible to achieve both efficiency and performance simultaneously? In this paper, we propose Mixture of Sparse Adapters (MoSA) as an affirmative answer to this question, enhancing adapter performance without introducing additional computational and storage overhead, which is shown in Figure 1(c). We start by splitting the standard adapter into several non-overlapping modules, each can be considered as a sparse expert. During training, to avoid extra computational costs, we use a non-routing stochastic activation mechanism. Each activated module uses a mask to filter gradients, achieving sparsity in parameter updates. We merge all sparse adapters into a complete dense adapter, achieving efficiency in both storage and computation during inference. Through this design, the potential of the original adapter is fully unleashed, maximizing the elimination of parameter redundancy.
@@
Rather than simply stacking two independent approaches, our design organically integrates the stochastically activated MoE with sparse training, where the two components mutually reinforce each other. On one hand, sparse update of parameters more efficiently utilizes training data, alleviating the data dilution caused by sparse activation of multiple experts. On the other hand, the training process with mixed experts ensures expressive capability and stability in downstream tasks, addressing the limitations of sparse training. To better facilitate the combination of the two components, we further propose a hierarchical sparse strategy. The dense down-projection layer provides robust intermediate features for the model, while multiple sparse up-projection layers increase the capacity of the model. It’s worth noting that our approach, as a general concept of enhancing parameter utilization, can be applied to various adapter structures and other addition-based PEFT methods. On 27 diverse downstream visual tasks, our MoSA consistently outperforms all other fine-tuning methods, including full fine-tuning. Compared to full fine-tuning, MoSA achieves a lead of 1.36% (on FGVC), 2.43% (on VTAB) and 1.51% (on GICD), while updating only around 1% of the backbone parameters. Compared to the standard adapter, MoSA achieves an improvement of 1.32% (on FGVC) and 1.06% (on VTAB) without adding any computational or storage overhead. Additionally, we conducted comprehensive ablation experiments to validate the effectiveness of each component in our design. The results demonstrate that MoSA is indeed an Adapter Tuning method that successfully balances efficiency and performance.
@@
  • We propose a novel Adapter Tuning method, namely MoSA, for fully unleashing the potential of the standard adapters. Through a mixed and sparse training approach involving splitting and merging, our method maximizes parameter efficiency, enhancing the performance of adapters in visual tasks.
  • MoSA best achieves a balance between efficiency and performance. It exhibits efficiency in all the sparsification, mixed training, and inference stages, and the mutual promotion between stochastic activation and sparse training further enhances performance.
  • We evaluate our method on a total of 27 downstream visual tasks spanning different domains, and MoSA significantly outperforms full fine-tuning as well as all other PEFT baselines, demonstrating the rationality of our design.
  • We conduct comprehensive ablation studies to explore various design choices, demonstrating the effectiveness of each component. Additionally, we showcase the consistent improvements brought by MoSA across multiple model scales, architectures, and different PEFT methods.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT). Recently, many large-scale pre-trained models (e.g. LLaMA2 [69] 70B, GPT-3 [5] 175B) have emerged in deep learning research, which can achieve excellent performance in a variety of downstream tasks. However, updating and storing all model parameters for each task has become far more expensive. PEFT achieves efficiency in training and storage by updating only a small fraction of parameters compared to pre-trained models [80, 46, 45, 36, 37, 47, 48, 50, 85, 27], among which Adapter Tuning [33, 64, 65, 8, 60, 61] is one of the most widely used. Due to space constraints, we focus exclusively on Adapter Tuning. For a broader overview of other PEFT methods, we recommend readers refer to surveys on tuning [14, 76]. Adapters [33, 64] are first introduced in natural language processing (NLP), achieving efficient knowledge transfer by only updating the parameters of newly added lightweight bottleneck modules. AdaptFormer [7] first applies adapters to visual recognition, achieving remarkable results in video understanding. Subsequent works [34, 39, 31] implement low-rank adaptation through various decomposition manners, further reducing the number of parameters required for fine-tuning. MoA [44] addresses the domain generalization through a MoE version of the aforementioned adaptation methods. Recently, SparseAdapter [30] enhances parameter efficiency through pruning of adapters before tuning. Although both MoE and sparsification can improve adapter performance, they also introduce additional computational costs. In this work, we propose a novel Adapter Tuning framework that enables two techniques to mutually enhance each other, ensuring both efficiency and performance.
@@
Mixture-of-Experts (MoE). The origin proposal of MoE [35] is targeted at enhancing model capacity. The earliest proposed technologies, known as soft MoE [55], integrate outputs of multiple experts through weighted summation. However, this technique significantly increases the computational cost during training. To address this issue, the sparsely-gated MoE [67, 84, 27] was introduced, selecting specific experts for activation and directly assigning data to them, thereby reducing the computational burden. Nonetheless, this method often resulted in load imbalance, with some experts becoming inactive in later training stages, risking system collapse. THOR [86] effectively mitigates this issue by implementing random activation, which not only enhances efficiency but also improves overall model performance. AdaMix [73] extends this mechanism to efficient tuning, treating each PEFT module as an expert and achieving performance improvements in NLP tasks. However, the stochastic activation of multiple experts may lead to the issue of data dilution issue, resulting in suboptimal performance when there is insufficient data for downstream tasks. To address this problem, we employ sparse training for each expert, reducing the data requirement per expert. This approach significantly enhances performance while maintaining the same computational efficiency as standard adapters.
@@
Pruning and sparse training. The primary goal of model pruning [26, 58, 54] aims to minimize deployment costs by reducing model parameters without significantly impacting its performance. Recent studies [24, 75, 1] have revealed that pruning can enhance model fine-tuning. A reduced number of trainable parameters can act as an additional regularization constraint, potentially improving performance [85]. As a PEFT method, the architecture of adapters [33, 64] inherently possesses few updatable parameters. SparseAdapter [1] extends this concept by pre-pruning the adapter, enhancing parameter efficiency while maintaining or even improving performance compared to standard adapters. However, excessive sparsity can lead to training instability and suboptimal results on certain datasets. We take one step further to solve aforementioned weakness via integrating the MoE paradigm into our sparse tuning process. This integration not only achieves a sparse network structure, but also implements sparse activation during the training phase, which overcomes the inherent constraints of using a single sparse network and increases the model capacity.
@@ 3 Method
Figure 2: Architecture design of Mixture of Sparse Adapters. In the training phase, MoSA stochastically activates a sparse adapter during each forward pass; in the inference phase, MoSA merges multiple sparse adapters into a complete one to enhance efficiency.
To achieve the optimal balance between efficiency and performance, we propose Mixture of Sparse Adapters (MoSA). The overall design of our method is shown in Figure 2. We first overview our method in Section 3.1. Subsequently, we describe how we split the standard adapter into multiple sparse adapters in Section 3.2. Then, we introduce the training strategy with stochastic activation of multiple sparse experts in Section 3.3. Finally, we merge the split sparse adapters into a complete one for efficiency during inference in Section 3.4.
@@ 3.1 Overview
where $x\in\mathbb{R}^{d}$ represents the input of the adapter, $f$ is the activation function, $\mathcal{W}^{\text{down}}\in\mathbb{R}^{d\times r}$ and $\mathcal{W}^{\text{up}}\in\mathbb{R}^{r\times d}$ denote the linear layers for down-projection and up-projection, respectively. $r$ is the bottleneck dimension, satisfying $r\ll d$, which allows for a reduction in the number of parameters in the adapter. Increasing $r$ can enhance the performance of the adapter but also increases the total number of parameters. SparseAdapter [30] further exploits the high efficiency of parameters by pruning $\mathcal{W}^{\text{down}}$ and $\mathcal{W}^{\text{up}}$ before tuning. In our work, we split the standard adapter into $N$ sparse adapter experts $\mathscr{W}^{\text{down}}=\{\mathcal{W}_{i}^{\text{down}}\}_{i=1}^{N},\mathscr{W}^{\text{up}}=\{\mathcal{W}_{i}^{\text{up}}\}_{i=1}^{N}$, and stochastically activate one of them during training for adaptation:
@@
where $i,j\in\{1,\cdots,N\}$, representing the stochastic selection of $N$ experts.
@@ 3.2 Sparse Adapter Splitting
To make the most of all the parameters within a single adapter, we first split the standard adapter into multiple sparse adapters, as illustrated in Figure 1(c). As all parameters in the adapter need to be updated, unlike [30], we do not adopt the task-specific pruning mechanism but instead employ a random splitting method to achieve as balanced grouping as possible, which also avoids the additional overhead of gradient computation based on downstream tasks before fine-tuning. Given the parameter of adapter $\mathcal{W}\in\mathbb{R}^{d\times r}$, we first generate an initial score $\mathcal{S}\sim\text{Uniform}(0,1)\in\mathbb{R}^{d\times r}$. Then, we split $\mathcal{W}$ into $N$ non-overlapping sparse adapters $\mathcal{W}_{i}:i\in\{1,\cdots,N\}$ by $\mathcal{S}$. Denoting the $i$-th $N$-quantile of all values in the matrix $\mathcal{S}$ as $s_{i}:i\in\{1,\cdots,N\}$, we obtain $N$ sparse masks $\mathcal{M}_{i}$ as follows:
@@
where $\varepsilon$ represents the update step size, and $\nabla\mathcal{L}(\mathcal{W}_{i})$ denotes the gradient of the task loss with respect to $\mathcal{W}_{i}$. In this way, the positions with a mask value of 0 have their gradients filtered to 0, thereby freezing the corresponding parameters and achieving sparse training for the adapter. Consequently, we obtain $N$ independent sparse adapters.